solve each of the following systems by sustitution.
8x-4y=16
y=2x-4

Answer by venugopalramana(3286)

solve each of the following systems by sustitution.
8x-4y=16....................I
y=2x-4....................II....PUT THIS IN EQN.I
8X-4(2X-4)=16
8X-8X+16=16
16=16.....OK...
HENCE THE EQNS.ARE DEPENDENT ANE THERE ARE INFINITE SETS OF SOLUTIONS..ANY SET WITH
X=ANY VALUE......SAY'0'
AND Y=2X-4=2*0-4=-4 IS A SOLUTION SET.
YOU CAN GIVE ANY VALUE TO X,FIND CORRESPONDING Y AND THAT IS THE SOLUTION SET.